OpenStudy (anonymous):

Solve this equation for r: (2+3)r – 6 = 31

OpenStudy (anonymous):

5-6=31

OpenStudy (mathmale):

Distributing that r would work, but to my mind it'd be far more efficient NOT to distribute the r. We want to isolate r, not have r in several terms.

OpenStudy (mathmale):

(2+3)r – 6 = 31 May I suggest adding 6 to both sides, to further isolate that r?

OpenStudy (anonymous):

im sorry im trying i never done these

OpenStudy (mathmale):

You're familiar with adding a number to both sides of an equation?

OpenStudy (anonymous):

add 6 to 5?

OpenStudy (mathmale):

Thank you for contributing, @Ludwig457.

OpenStudy (mathmale):

What I see is (2+3)r-6=31. If you combine 2 and 3, you get 2+3 = 5. Thus, (2+3)r=5r, and 5r-6 = 31. +6= +6 ------- --- Please complete this work.

OpenStudy (anonymous):

By inspection: r = (31+6)/5

OpenStudy (mathmale):

Couldn't agree more! But I sense that @LaBlueEyes2013 would appreciate a step-by-step approach the first time through a problem such as this one.

OpenStudy (mathmale):

(2+3)r – 6 = 31 1. combine 2 and 3: 2+3 = 5, and so (2+3)r = 5r 2. rewrite the equation as 5r - 6 = 31. 3. Add 6 to both sides and simplify the resulting equation.